OpenStudy (anonymous):

Need hypothesis testing help!!! A real estate agent is working for a developer who claims that the average commute time to downtown is 20 minutes with a standard deviation of 7 minutes. Stephon is an independent real estate agent and wants to check the times for his client. He took a random sample of 15 commute times and found an average of 26 minutes. He did hypothesis testing using a significance level of 5%. Which conclusion could he make? The z-statistic is 0.22, so the null hypothesis should not be rejected. The z-statistic is 1.69, so the null hypothesis should not be rejected. The z-statistic is 3.32, so the null hypothesis should be rejected. The z-statistic is 3.67, so the null hypothesis should be rejected.

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

@jrain35 Do you know how to calculate the z-statistic?

OpenStudy (anonymous):

no thats my problem with these questions :/

OpenStudy (kropot72):

What learning resources have been made available to you?

OpenStudy (anonymous):

none that i know of

OpenStudy (kropot72):

What course are you taking?

OpenStudy (anonymous):

Algebra 2

OpenStudy (kropot72):

The formula for finding the z-statistic is as follows: \[\large Z=\frac{\bar {X}-\mu}{\frac{\sigma}{\sqrt{n}}}\]

OpenStudy (kropot72):

Plugging in the given values, we get: \[\large Z=\frac{26-20}{\frac{7}{\sqrt{15}}}=you\ can\ calculate\]

OpenStudy (anonymous):

so x is sample size?

OpenStudy (anonymous):

or the mean?

OpenStudy (anonymous):

I think its the mean

OpenStudy (kropot72):

X-bar is the sample mean. Mu is the proposed value for the mean in the null hypothesis. Sigma is the standard deviation. n is the sample size.

OpenStudy (anonymous):

alright thanks a ton

OpenStudy (kropot72):

You're welcome :)